12.755g of a metal initially at 99.0C are placed into 75.0g of water initially at 2.30C. The final temperature of the water is 24.5C. Calculate the specific heat of the metal. The specific heat of water is 4.184J/g.C



1
Expert's answer
2021-09-17T02:08:44-0400

∆T = 99-24.5 = 44.5°C

Mass = 12.755

= 12.755 × 44.5 × 4.184 = 2374.83

= 2.2375kJ/g.c